OpenStudy (anonymous):

Please help... will fan and medal :) For the given statement Pn, write the statements P1, Pk, and Pk+1. (2 points) 2 + 4 + 6 + . . . + 2n = n(n+1)

OpenStudy (amistre64):

whats the issue?

OpenStudy (anonymous):

im not sure if im going to do the problem right and just need assistance

OpenStudy (amistre64):

well, show me what you think should happen and ill correct as needed

OpenStudy (anonymous):

P1 = 1(1+1)

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Pk= k(k+1)

OpenStudy (anonymous):

Pk+1 = k+1(k+1+1) ??

OpenStudy (amistre64):

thats only partly right P1 = 2(1) = 1(1+1)

OpenStudy (amistre64):

Pk is just Pn rewritten with k instead of n

OpenStudy (anonymous):

Yeah i know that much im just not sure if i was doing it right for this specific problem.. i think im over thinking it ?

OpenStudy (amistre64):

Pk = 2 + 4 + 6 + . . . + 2k = k(k+1)

OpenStudy (amistre64):

P(k+1) = 2 + 4 + 6 + . . . + 2k + 2(k+1) = k(k+1) + 2(k+1) | addon| | addon|

OpenStudy (amistre64):

it looks like you might be starting proofs by induction?

OpenStudy (anonymous):

yeah ..

OpenStudy (anonymous):

Am i on the right track though ?

OpenStudy (amistre64):

then it important to understand that: P(k+1) = P(k) + (addon)

OpenStudy (amistre64):

your dont fine so far, the idea is that we want to convert P(k)+(addon) into the same form as P(k) in other words, we know the form of P(k) is good therefore its the form we want to establish P(k) = k(k+1) P(k+1) = P(k) + (addon) = k(k+1)+ (addon) and then algebra the right side to the form (k+1)(k+1+1)

OpenStudy (amistre64):

*** doing fine ....

OpenStudy (anonymous):

idk what addon is :x

OpenStudy (amistre64):

its the 'rule' for the end term in P(n) in this case if 2(k+1) since the nth term is defined as 2n

OpenStudy (amistre64):

2 + 4 + 6 + . . . + 2n ^^^ rule for the nth term P(k) = 2 + 4 + 6 + . . . + 2k P(k+1) = [2 + 4 + 6 + . . . + 2k] + [2(k+1)] P(k+1) = P(k) + 2(k+1)

OpenStudy (anonymous):

And thats it then?

OpenStudy (amistre64):

thats what its going for yes now remember that what you add to one side of an equals sign, you also add to the other \[P(k)= 2 + 4 + 6 + . . . + 2k = k(k+1)\] \[P(k+1)= \underbrace{\color{red}{2 + 4 + 6 + . . . + 2k}}_{P(k)}+2(k+1) = \underbrace{\color{red}{k(k+1)}}_{P(k)}+2(k+1)\]

OpenStudy (amistre64):

we can now factor the right side since (k+1) is common to both terms \[k(\color{green}{k+1})+2(\color{green}{k+1})\implies (\color{green}{k+1})(k+2)\]

OpenStudy (anonymous):

I think im confusing myself lol im sorry ..

OpenStudy (amistre64):

well,this is just a preview of the thought process you have to work with when doing induction proofs. just trying to get you exposed to it now so that its not such a shock in a few days :)

OpenStudy (anonymous):

Oh .. im doing online schooling and this is the last question of this assignment .. and idk this one i just got stumped on it .. its from a previous lesson

OpenStudy (anonymous):

Just to be clear.. this is not correct? So for the answer i would then write : p1= 2(1)=1(1+1) Pk= 2k=k(k+1) pK+1= 2(k+1)=k+1(k+1+1)

OpenStudy (amistre64):

\[P(1):~2(1)=1(1+1)\] \[P(k):~2+4+6+...+2k=k(k+1)\] \[P(k+1):~2+4+6+...+2k+2(k+1)=k(k+1)+2(k+1)\\ \hspace{4em}\color{red}{\implies}(k+1)(k+1+1)\]

OpenStudy (amistre64):

P(n) is not defined to be the last term .. its the sum of all the terms

OpenStudy (anonymous):

gotchaaa .. thank you for that (:

OpenStudy (amistre64):

P(1) = 2(1) P(2) = 2(1) + 2(2) P(3) = 2(1) + 2(2) + 2(3) P(n) = 2(1) + 2(2) + 2(3) + ... + 2(n) P(k) is not 2k

OpenStudy (amistre64):

good :)

OpenStudy (anonymous):

OHHHHHHHH okay .. that makes since.. i didnt realize thats how i was writing it.. thank you so much again

OpenStudy (anonymous):

sense*